OpenStudy (anonymous):

What is the height of a rectangular based pyramid with a volume of three hundred cubic meters and base dimensions of fifteen meters and ten meters?

OpenStudy (anonymous):

1) Vpyramid = 1/3 * Vcube = 27m^3 Ans. B 2) 3.14 * r^2 * h is the volume of the cylinder , 1/3 of that is the volume of the cone Ans. B 3) h = V/bw = 300 / (15 * 10) = 2m Ans A 4) rhombus = 9 * 4 = 36 triangle = 1/2 * 12 * 4 = 24 circle = pi * 2^2 = 4 * pi ~= 14 trapezoid = 1/2 * (3 + 11) * 6 = 42 Ans D.

OpenStudy (anonymous):

thanks but i still don't really understand

OpenStudy (anonymous):

900 meters cubed

OpenStudy (anonymous):

do you want me to send you my choices

OpenStudy (anonymous):

2 meter 6 meters eighteen meters fifty-four meters

OpenStudy (anonymous):

the answer is 2 meters